The perimeter of a triangle is 45 centimeters. The first side is 3 centimeters shorter than the second side. The third side is 6 centimeters shorter than twice the length of the first side. Find the length of the shortest side.
Answer by hamsanash1981@gmail.com(151)

Perimeter = 45
Let the second side be x, then first side = x-3, and third side = 2(x-3)-6 =2x-12
first side +second side + third side = 45
x-3+x+2x-12 = 45
4x -15 = 45
4x = 60
x = 15
first side = 15-3 = 12
second side = 15
third side = 2(12)-6 =18
The length of the shorter side is 12.